### v1

# XAI4LLM. Let Machine Learning Models and LLMs Collaborate for Enhanced In-Context Learning in Healthcare

Abstract.
The integration of Large Language Models (LLMs) into healthcare diagnostics offers a promising avenue for clinical decision-making. This study outlines the development of a novel method for zero-shot/few-shot in-context learning (ICL) by integrating medical domain knowledge using a multi-layered structured prompt. We also explore the efficacy of two communication styles between the user and LLMs: the Numerical Conversational (NC) style, which processes data incrementally, and the Natural Language Single-Turn (NL-ST) style, which employs long narrative prompts.
Our study systematically evaluates the diagnostic accuracy and risk factors, including gender bias and false negative rates, using a dataset of 920 patient records in various few-shot scenarios. Results indicate that traditional clinical machine learning (ML) models generally outperform LLMs in zero-shot and few-shot settings. However, the performance gap narrows significantly when employing few-shot examples alongside effective explainable AI (XAI) methods as sources of domain knowledge. Moreover, with sufficient time and an increased number of examples, the conversational style (NC) nearly matches the performance of ML models. Most notably, LLMs demonstrate comparable or superior cost-sensitive accuracy relative to ML models.
This research confirms that, with appropriate domain knowledge and tailored communication strategies, LLMs can significantly enhance diagnostic processes. The findings highlight the importance of optimizing the number of training examples and communication styles to improve accuracy and reduce biases in LLM applications.
Large language models, ChatGPT, Clinical decision-support, XAI, Generative, Prompt, Conversation
copyright: acmlicensedjournalyear: 2024doi: XXXXXXX.XXXXXXX
1. Introduction
Context. Unlike traditional machine-learning models, which are often limited to functioning as “narrow-expert” systems optimized for specific tasks, modern generative models demonstrate a remarkable ability to learn from and sample a broad range of data distributions, for example, those obtained from vast, diverse sources from large-scale internet, encompassing text and visual content. This capability enables them to generate authentic-looking data that closely resembles the training data. Furthermore, recently, the advancement of natural language processing (NLP), combined with increased computational power and data availability, has spurred the development of large-scale generative models, known as large language models (LLMs), such as ChatGPT (OpenAI, 2022), and Gemini (Team et al., 2023). This new form of learning does not rely on limited, narrowly defined datasets to support a traditional $train\rightarrow prediction\rightarrow label$ paradigm (in a supervised setting). Instead, pretrained large language models (pLMS) leverage massive parameter data to facilitate a direct $prediction\rightarrow label$ process, potentially conserving resources and enhancing nuanced learning. Consequently, once provided with the appropriate context and domain knowledge, these models can be optimized for use in various downstream tasks. Overall, thanks to their wide-ranging, generalist emergent capabilities (Schaeffer et al., 2023), these models have introduced a myriad of opportunities for research across various AI disciplines (Imani et al., 2023; Zhang et al., 2023; Jiang et al., 2023; Xu et al., 2023; Li et al., 2023).
PLMS offer significant opportunities and benefits in critical domains such as the healthcare sector, by providing critical support across various tasks such as medical assessment, question answering, and patient classification (cf. Section 2.2). They hold the premise to enhance the accuracy of diagnostics, improve patient interactions, and integrate diverse data types into clinical decision-making, thereby impacting global healthcare delivery. This paper explores the use of LLMs for zero/few-shot in-context learning (ICL) in healthcare setting, focusing specifically on “heart risk prediction”. By employing the general-purpose LLM, ChatGPT, as a case study, we aim to demonstrate how these models can improve diagnostic accuracy in healthcare decision-making. Although our findings concentrate on heart risk prediction, the methodologies and insights are applicable to a broad range of health-related tasks, and LLM types.
Table 1. Comprehensive Research Challenges in Healthcare Decision Making with LLMs
Research Challenge and Description
Context Integration. The accuracy of LLM outputs in healthcare decision making (DM) relies on integrating “domain-specific knowledge.” Given that pretrained models such as ChatGPT are designed as general-purpose tools, it becomes imperative to question, “is a general-purpose LLM adequate for clinical DM, and how can it be optimized to enhance its domain awareness?” For example, integrating diagnostic guidelines from medical databases can help tailor LLM responses to specific clinical questions. Identifying optimal methods for incorporating context from human experts with medical knowledge or external data sources into LLM-based DM systems is a key research area.
Numerical vs. Natural Language Profile. This aspect addresses the challenge of optimally constructing patient profiles within prompts for enhanced in-context learning (ICL). Medical diagnostics depend heavily on quantitative data-lab results, vital signs, and more—requiring precise interpretation. The challenge lies in enhancing the abilities of LLMs to accurately process and understand these numerical values within a medical context, “How can we bridge the gap between numerical and linguistic proficiency in LLMs for healthcare DM?”
Communication Styles with LLMs. We recognize in this work, the impact of multi-turn (conversational) versus single-turn interactions on LLMs’ diagnostic precision. Communication styles with LLMs leverage structured interaction patterns, employing “<User>” and “<System>” tags to facilitate conversations between the user and the LLM. Two predominant styles are recognized: multi-turn (“conversational”) and single-turn interactions. The conversational style utilizes a sequence of short, interactive exchanges leading to a heart disease risk assessment, enabling incremental information gathering similar to the approach of a clinician approach with a patient, while the latter compiles all patient details into one (or few) comprehensive prompts. The challenge revolves around determining which communication style — the dynamic, detail-oriented conversational or the direct, comprehensive single-turn — effectively enhances diagnostic precision of the LLMs.
Reasoning Mode and the Interplay with Communication Style. This step explores the effectiveness of various LLM reasoning modes, with a particular focus on the chain-of-thought (COT), in enhancing interpretability and accuracy within complex diagnostics. We aim to understand whether COT inherently benefits the conversational data already presented to the system, or if it is more effectively utilized within extensive, single-turn prompts. Overall, the dynamic interplay between an LLM reasoning style and the user-system interaction emerges as a compelling area for advancing the ability of LLMs to generate more precise and insightful responses, a nuanced aspect studied in this work.
Risk Evaluation in Clinical Decision-Making. While enhancing the accuracy of LLMs in the context of clinical DM is essential, and their performance comparison with classical ML models is intriguing, an equally important aspect is understanding the inherent risks associated with deploying these systems in clinical settings. Among the various risks, we specifically focus on two: false negatives and biases (Panigutti et al., 2021; Amigó et al., 2023; Makhlouf et al., 2021). False negatives are given particular emphasis in our research due to the grave implications of missing a risk, which can be more critical than incorrectly identifying a non-existent condition (false positives). Another risk that we aim to deepen our understanding of is gender bias, which could exacerbate existing disparities in healthcare outcomes. While we do not propose specific mitigation strategies within the scope of this study, recognizing and highlighting these risks is crucial for future research and development efforts aimed at making LLM-based diagnostic systems safer and more equitable for clinical decision-making.
Challenge. Integrating Large Language Models (LLMs) into healthcare decision-making (DM) is inherently a complex task. It requires an understanding of and addressing key issues to fully leverage their potential. This work identifies critical areas for the integration of LLMs in clinical DM, which include: (i) enriching general-purpose LLMs with the specificities of the task at hand by integrating medical context – domain knowledge; (ii) bridging the gap between numerical datasets and the natural language processing capabilities of LLMs; (iii) identifying optimal communication styles between users and systems for gathering better or more precise response from LLMs –a nuanced aspect identified in our research; (iv) assessing the impact of different reasoning modes (e.g., direct request vs. chain-of-thought) on diagnostic clarity; and (v) evaluating LLMs in clinical DM not only for better accuracy but also for unveiling the associated risks.
Addressing these challenges is important for the effective use of LLMs in healthcare DM. For a detailed discussion of each challenge, we have created Table 1, that summarizes each challenge in more detail. Please note that some of the challenges might be deemed domain-independent; however, for the purpose of this work, we discuss and illustrate them in the healthcare setting.
1.1. Research Objectives and Contributions
Our primary objective is to investigate the applicability of LLMs in healthcare decision-making, focusing specifically on zero/few-shot ICL for heart risk prediction. We aim to enhance diagnostic accuracy by integrating domain-specific medical knowledge into the decision-making process facilitated by LLMs. Thus, this research seeks to fill the gap in current knowledge by presenting a detailed examination of LLM application in medical contexts, contrasting it with traditional ML approaches. This work introduces the following main contributions:
  1. (1)
    Systematic Structured Prompt Construction. We propose a systematic approach to ICL within the medical domain by designing a “multi-layered structured template,” aimed at seamlessly integrating several elements related to the clinical task at hand. Most importantly, it incorporates medical domain knowledge into the proposed LLM-based Decision-Making (DM) system. Specifically, the proposed prompt structure in our work addresses these nuanced areas for research:
    • (i)
      User profile generation, handling two types of representation: numerical and natural language (NL).
    • (ii)
      Adding domain knowledge to enrich the prompts with medical knowledge. Building on our previous work (Nazary et al., 2023), the system sources this domain knowledge from trained Machine Learning (ML) models through feature-level explanations. This approach mirrors the insights one might obtain from doctors or medical literature, emphasizing the value of such domain-specific knowledge.
  2. (2)
    Optimizing Communication Styles and Profiles in LLM-based Diagnostics. Our research recognizes two communication styles, intricately combined with patient profile construction: (i) the “conversational” style, which employs multi-turn interactions utilizing numerical data to mimic the dynamic, incremental nature of clinician-patient dialogues, and (ii) the “Single-Long Prompt” style, which integrates all patient details into a comprehensive narrative for a holistic assessment similar to a review of clinicians of a complete medical file before diagnosis. This approach is motivated by the need to understand how the structured presentation of patient data—whether through a detailed, step-by-step conversation or a broad, single overview—can effectively harness the capabilities of LLM to yield accurate heart disease risk assessments. Examples of these two approaches are shown in the following section.
Illustrative Example. In Figure 1, we illustrate the essence of our contributions through a visual comparison of two distinct communication styles within the context of LLM-based diagnostics. The left column showcases the “Conversational” style, where the interaction unfolds through a series of multi-turn exchanges. This method mimics the incremental and dynamic nature of clinician-patient conversations, gradually compiling a comprehensive user profile that blends both numerical data and natural language descriptions. It is designed to engage the LLM in a way that closely parallels the iterative diagnostic process in clinical settings. Conversely, the right column demonstrates the “Single-Turn Prompt” approach, which compiles all relevant patient information into one extensive narrative. This style aims to provide the LLM with a complete, integrative overview of situation of the patient in one round, similar to a clinician reviewing the entire medical history of a patient before making a diagnosis. The contrast between these columns highlights our systematic structured prompt construction ability to adapt the presentation of patient data to optimize LLM-based diagnostic accuracy.
{mdframed}
Key Insight and Assumption: Overall, noting that the ML models have access to in-domain knowledge, compared to pre-trained Language Learning Models (pLLMs), which are: (i) general-purpose but trained on massive datasets, and (ii) rely on limited few-shot examples in few-shot ICL settings, we posit that insights obtained from ML for the task at hand, can better refine those of pLMMs. Our overarching insight is that by allowing ML models and LLMs to collaborate, we can more effectively harness their strengths for enhanced DM in a complementary manner. Ultimately, Explainable Artificial Intelligence (XAI) serves as the main catalyst to integrate in-domain knowledge with general-purpose LLMs, thus bridging the gap between data-driven insights and actionable understanding.
Figure 1. Comparison of two communication syles with LLMs as introduced in this research
2. Related work
2.1. Generative models and LLMs
Generative models have been a cornerstone of machine learning (ML) for some time, showcasing notable success across various applications within the field. Their fundamental concept revolves around understanding the underlying distribution of the data they are trained on—and sampling from it—in order to generate new, similar data. This can range from text and images to more complex outputs. Numerous architectures have been developed, each with its own unique method for learning and reproducing data distributions (Harshvardhan et al., 2020). These include GANs [^1] Generative Adversarial Networks, VAEs [^2] Variational Autoencoders, Diffusion Models [^3] A class of generative models that simulate a diffusion process to gradually transform noise into data, Conditional Generative models [^4] Models that generate data conditioned on certain inputs etc.
Recently, especially after 2023, the popularity of these models has experienced a significant surge, reminiscent of the rise in deep neural networks around 2015 following the introduction of Convolutional Neural Networks (CNNs) (Li et al., 2021). This increase in interest is largely due to the advent of large pre-trained language models (LLMs), such as ChatGPT, along with numerous successors such as LLaMA, Claude, Cohere, LaMDA, PaLM, and GPT-Neo, among others (Zhao et al., 2023; Xi et al., 2023), which have collectively pushed the boundaries of what pre-trained LLMs can achieve in terms of understanding and generating human-like text.
In the context of LLMs, we can recognize two main categories of generative models, each exemplified within healthcare domain.
  1. (1)
    Directly Trained Models. These models are explicitly trained on specific datasets, such as patient symptoms, test results, and health outcomes, to learn and sample from the distribution of patient data (Torfi and Fox, 2020; Choi et al., 2017). This approach benefits from closely aligning the model with the unique characteristics of individual health datasets, enabling predictions of patient outcomes with a detailed comprehension of their medical histories.
  2. (2)
    Pretrained Generative Models. These models utilize the vast knowledge acquired from being trained on a broad spectrum of data, textual, visual (and possibly tabular). By fine-tuning these models on disease-specific or demographic-specific datasets, their capacity to accurately interpret complex medical scenarios and patient histories is enhanced, thus improving their ability to offer precise diagnostic insights or personalized treatment plans. They can be used in the following scenarios (Ge et al., 2023; Moor et al., 2023; Spolaôr et al., 2024; Henriksson et al., 2023; Shentu and Al Moubayed, 2024):
    1. a
      Zero- and Few-shot In-Context Learning. This represents a remarkable emergent capability of LLMs, showcasing exceptional zero- and few-shot ICL. For example, these techniques enable machine learning models to apply previously learned concepts to novel problems without the need for extensive further training. This is particularly important in emerging health issues, where data may be limited;
    2. b
      Fine-Tuning. This process involves adjusting pre-trained models with in-domain data, such as targeted medical datasets, thereby improving their effectiveness for specific health conditions and enhancing the personalization of care.
    3. c
      Embedding Generation for Downstream Training. This technique is particularly useful for interpreting complex medical data such as electronic health records or imaging data, where embeddings can capture intricate patterns and relationships (Yang et al., 2023).
As the role of AI/ML in critical domains such as healthcare continues to mature, it is important to understand the challenges of integrating these new technologies into clinical practice, and understand their associated risks. Our work at hand is focused on the first category here, zero- and few-shot ICL.
2.2. Overview of LLMs in the healthcare
Large Language Models (LLMs) and, more broadly, generative AI, hold the potential for widespread application across various domands and task, including finance (Wu et al., 2023; Deldjoo, 2023), healthcare (Nazary et al., 2023; Peng et al., 2023; Ghosh et al., 2024), agriculture (Rezayi et al., 2023), recommender systems (Deldjoo, 2024; Deldjoo et al., 2024) and more, as explored in recent research works. Of particular importance are critical sectors such as healthcare and finance, where the societal impact is significant. In these areas, it is crucial not only to focus on the positives and benefits offered by the system in terms of better predictions but also to critically understand and evaluate the associated risks and challenges.
The healthcare sector is one of the most critical areas where LLMs or more broadly generative AI could have a transformative effect, but it also poses risks since decisions directly impact human lives. On the positive side, a number of recent surveys (Yang et al., 2023; Nazi and Peng, 2023; Chen and Esmaeilzadeh, 2024; Umer and Adnan, 2024) provide a good frame of refrence on various tasks that LLMs can achieve within the healthcare sector. These include medical assessment, medical question answering, scientific writing, eHealth care, and patient classification.
For instance, the survey by Yang et al. (2023) examines the potential applications of LLMs in healthcare, highlighting their role in developing domain-specific models to enhance clinical utility. These models aim to minimize errors and maximize effectiveness for global healthcare delivery, either by building from scratch or refining existing larger models. This approach facilitates decentralized research, enabling researchers worldwide to tailor LLM applications to specific clinical tasks. Moreover, the authors detail how LLMs integrate into the healthcare process, which is categorized into three main stages: (1) Pre-consultation, involving screening and initial medical consultations for symptomatic or asymptomatic patients; (2) Diagnosis, which includes collecting patient history, conducting physical examinations, and performing diagnostic investigations such as CT scans; and (3) Management, encompassing the prescription of medications, patient education, counseling, and management of insurance claims. The approach of examining LLM applications according to timelines and temporal screening presents an interesting method for categorizing these interventions.
Another notable survey by Nazi and Peng (2023) presents an alternative interesting categorization of the various applications of LLMs in healthcare, organized for different stakeholders: for (1) patients, LLMs support health education, lifestyle advice, patient care, and function through virtual medical assistants and health chatbots; for (2) doctors, they extend to treatment planning, collaborative diagnosis, medical literature analysis, and clinical decision support; and for (3) central healthcare applications, they include drug discovery, automated medical report synthesis, and radiology and imaging. This structured approach underscores the wide-ranging impact of LLMs in enhancing and personalizing healthcare services across different levels of medical practice.
2.3. LLMs for Automated Healthcare Diagnosis
The integration of LLMs into healthcare has been an area of significant research interest, aiming to enhance diagnostic accuracy and decision-making processes. The work at hand contributes to advancing the application of LLMs in medical diagnostics, each adopting unique approaches to harness the capabilities of LLMs in addressing complex healthcare challenges. A few recent research works are reviewed in the following:
The research work by Sayin et al. (2024) emphasize the role of LLMs in supporting medical decision-making but focuses on correcting physician errors during diagnostic processes. This research highlights the importance of prompt design in the performance of LLMs, similar to our approach. However it has a specific focus on interactive scenarios between physicians and LLMs, aimed at improving diagnostic outcomes based on physician-LLM collaboration. Kusa et al. (2023) explore the impact of user beliefs and prompt formulation on health diagnoses, demonstrating the sensitivity of LLMs to input variations. This study complements our findings by addressing the challenges in ensuring consistency in LLM-generated diagnoses, emphasizing one the role of tailored prompt design to mitigate the impact of user input biases. Galitsky (2024) studies the personalization in LLM applications, particularly in generating tailored medical recommendations. While our focus is on diagnostic accuracy, the work by Gaitsky et al. underscores the potential of LLMs in personalizing healthcare interactions, which could enhance the user-specific applicability of diagnostic tools. Gramopadhye et al. (2024) explore chain-of-thought (CoT) reasoning in LLMs for medical question answering, a technique that is also used in our work together with the structured prompt design by facilitating detailed, step-by-step reasoning processes in LLMs. This approach aims at improving the interpretability and accuracy of LLM responses, which aligns with our goal to enhance diagnostic precision through detailed data analysis and reasoning.
Positioning our Work. The collective insights from these papers underscore the diverse applications of LLMs in healthcare—from improving diagnostic accuracy to personalizing patient interactions and correcting physician errors. Our contribution to this body of research involves a structured method for integrating comprehensive medical knowledge into LLMs. By systematically organizing patient data and medical knowledge into structured prompts and feeding them through conversational and narrative profiles, LLMs can achieve higher diagnostic accuracy and reliability. This not only supports clinical decision-making but also opens avenues for future research into optimizing LLM frameworks for various medical applications, ensuring that these systems are effectively tailored to meet specific clinical needs.
The integration of LLMs into healthcare diagnostics is an area of significant research interest, aiming to enhance both diagnostic accuracy and decision-making processes. Our paper contributes to advancing the application of LLMs in medical diagnostics by proposing a systematic approach that involves structured prompts integrating medical knowledge into the decision-making process, thus enhancing the capabilities of LLMs in analyzing patient data. In a previous version of this work (Nazary et al., 2023), we leveraged ChatGPT alongside few-shot and zero-shot ICL to enhance decision-making in healthcare. This approach focused on extracting and incorporating domain knowledge from trained ML models into prompt templates for healthcare applications, aiming to harness the adaptability of ChatGPT for specific challenges in healthcare decision support. The current work at hand can be seen as an extension of our previous work but does this in a much more systematic way, and explores various unlooked or overlooked areas, such as optimized integration of domain-specific knowledge into prompts, and the design of systematic structured prompt templates, optimizing communication styles between LLMs and users for enhanced diagnostic accuracy. It also addresses the challenge of processing and interpreting numerical data within medical contexts, and evaluating the inherent risks such as false negatives and biases in LLM-based clinical decision-making systems.
3. Foundations
This section will briefly discuss the foundations of this work, contrasting generative versus discriminative models (Section 3.1), and connecting this to the approach of LLMs (Section 3.2). It would basically help us understand the application of LLMs using prompt-based ICL for binary classification.
3.1. Generative and Discriminative Models
Generative and Discriminative models are two main methods in pattern recognition and machine learning tasks. Their fundamental difference lies in their approach to modeling and application:
  • Discriminative Models: These models focus on the probability of a label ($y$) given an observable (features $x$), formalized as:
    (1) $P(Y|X=x)$
    These models are widely used for prediction tasks, such as classification or regression. Discriminative models estimate the likelihood of a label based on observed features, effectively discriminating between different kinds of data instances.
  • Generative Models: These models aim to model the distribution of observables within a dataset $\mathcal{X}$ given a label $y$. This is formally represented as:
    (2) $P(X|Y=y)$
    Such models are thus proper for generating new data instances by learning the joint probability distribution $P(X,Y)$ or simply $P(X)$ when labels are not available. For instance, after learning the general characteristics of cats from a dataset, a generative model can create a new image that captures the essence of a cat, as it understands the joint probability distribution of the data and labels, $P(X,Y)$. This capability can be expressed as:
[…]
    Here, $X_{\text{new}}$ represents a new instance (e.g., an image of a cat) generated by the model, drawing from its learned distribution of features associated with cats. Thus, generative models can not only understand the characteristics that define each class but also produce new examples that reflect those insights.
3.2. Utilizing pLLMs in Discriminative Tasks via ICL
This section aims to bridge the conceptual gap between the probabilistic approaches of discriminative and generative models and their application within LLMs, especially in the context of binary classification tasks.
Given a dataset $\{(x_{i},y_{i})\}_{i=1}^{N}$ where $x_{i}\in\mathcal{X}$ represents the input features and $y_{i}\in\{0,1\}$ denotes the binary labels, the primary goal of a learning model is to accurately predict the label of unseen data points through a learned mapping $f$. This process typically involves minimizing a loss function, such as cross-entropy loss, to reduce the discrepancy between the predicted and actual labels.
3.2.1. Discriminative approach to binary classification
Discriminative models, or direct learning models, seek to map input features $X$ to labels $Y$ via a function $f$:
(4) $f:X\rightarrow Y$
In this approach, the objective of the models –often referred to as a learning model– is to learn the mapping $f$ to accurately predict the label of unseen data points. This is achieved by minimizing a loss function, such as cross-entropy, which quantifies the discrepancy between the predicted labels and the actual labels.
3.2.2. Generative pLLMs’ approach to binary classification
In contrast, generative models, herein referred to as “foundation models,” employ a distinct tactic. Instead of directly mapping inputs to labels, they aim to comprehend and replicate the data distribution to generate new, plausible data points. For binary classification, these models are adapted through prompt-based techniques to function as classifiers:
(5) $G:(prompt(x),\theta)\rightarrow y$
Here, $prompt(x)$ is a mechanism that reformats the input $x$ into a configuration amenable to the foundation model $G$, integrating the binary classification task within the operational framework of the model. Parameterized by $\theta$, the model subsequently produces a response construed as the class label $y$. Thus, this approach leverages the generative capabilities of the model to perform classification, relying on the internal understanding of the model from the task as framed by the prompt.
4. Proposed XAI4LLM Framework
This section introduces the architecture of our XAI4LLM framework tailored for clinical settings, aiming to utilize the vast capabilities of pretrained Large Language Models (pLLMs) endowed with billion-scale parameters. This strategy seeks to improve the accuracy of clinical DM. Figure 2 outlines the main components of the proposed XAI4LLM framework, structured around two main layers: prompt construction (cf. Section 4.1) and LLM interaction and prediction layer (cf. Section 4.2). Although our system is specifically devised for binary classification tasks (heart risk assessment), employing ChatGPT as a popular example of pLLMs, the methodologies, and insights gained from this work have the potential to be extended to other types of pLLMs, and to address more complex tasks beyond binary classification.
4.1. Prompt construction layer.
The primary objective of this layer is to convert numerical data from the test dataset into a format comprehensible by LLMs, thereby ensuring the patient data is contextualized appropriately. This entails incorporating a detailed task description within the prompt and integrating domain knowledge gleaned from training a suite of ML models. These components are subsequently woven into a structured multi-layer prompt template, which formats the textual data suitably for LLM interpretation. The main elements of the “prompt construction layer” consist of:
  • a.
    User profile generation. (cf. Section 4.1.1)
  • b.
    Domain knowledge enrichment (cf. Section 4.1.2)
  • c.
    Few-shot learning (cf. Section 4.1.3)
  • d.
    Infusing data within a structured prompt template (cf. Section 4.1.4)
Figure 2. Flowchart illustrating the conceptual structure of our proposed LLM4Health Framework
4.1.1. User Profile Generation.
Working with classical ML problems in the health domain often involves dealing with tabular data, typically in the form of numeric values (e.g., laboratory test results, vital signs). Given that LLMs are inherently designed to process and understand textual input, a methodological approach is required to convert numerical patient data into a format that LLMs can effectively engage with. The prompt construction layer is dedicated to this purpose, aiming to transform patient medical information and lab results (for instance, white blood cell count, glucose levels, blood pressure readings) into a textual format that LLMs can interpret. This transformation facilitates the comprehension and response capabilities of LLMs, bridging the inherent gap between numerical data and textual queries. In this context, we examine the efficiency of two distinct methodologies for adapting patient data into a textual narrative. This is applied both for the target patient in question and few-shot examples.
  1. (1)
    Numerical representation. In this method, we retain the numerical representation of patient features, reflecting the raw format of data. This style aligns with traditional ML input methods, allowing the LLM to process quantifiable data directly.
    Example:
        Age: 54, Sex: 1.0, CP: 3.0, Trestbps: 150, ...
  2. (2)
    Natural language (NL) representation Contrasting the conventional approach, the narrative style converts patient data into a comprehensive, single-turn long prompt. This method thus creates a story-like description that integrates the patient clinical features into a coherent narrative, aiming to leverage/test the language understanding capabilities of LLMs:
    Example:
    “A 54-year-old male presenting with non-anginal pain, a blood pressure reading of 150 mmHg, and a cholesterol level of 223 mg/dl, …”
4.1.2. Domain knowledge enrichment
In our enhancement of the XAI4LLM framework, we have enhanced our domain knowledge enrichment process, building upon the work outlined in our previous research (Nazary et al., 2023). The previous approach supplied the ML model with a straightforward explanation of feature-level attributes, essentially providing a list of the most important features. We have improved this step to categorize features into tiers of significance—high, medium, and low importance—leveraging insights from feature importance analysis derived from ML models. This method entails analyzing the distribution of feature importance scores and employing statistical quantiles to define the boundaries between categories.
Example:
Feature Importance Analysis:
  • Key Factors: cp, chol, oldpeak, thalach
  • Additional Factors: exang, trestbps, thal
As an example, the scatter plot in Figure 3 visualizes the importance of various features in a predictive model, mapped along quantile boundaries. The vertical axis lists the features, while the horizontal axis measures their importance. Color-coded dots represent the importance levels: green for high importance, blue for moderate, and light blue for low or neutral importance. Dashed lines denote the 33rd and 67th percentiles, categorizing the features into tiers based on their impact on the output of the model, related to heart disease risk assessment.
Figure 3. Scatter plot of feature importance with quantile boundaries, depicting the varying influence of XGBmethod on heart disease risk assessment. The color coding—green for high importance, blue for moderate, and light blue for low importance—alongside percentile markers, provides a clear hierarchical visualization of the factors contributing to the prediction of the model.
4.1.3. Few-Shot Learning.
In few-shot learning mode, this step supplies illustrative examples from the training dataset, illustrating diverse patient assessments. Each example, chosen at random, aims to highlight various risk factors and their impact on the outcome of heart disease risk assessments. These examples can be highly helpful for the learning capability of the model but also showcase the multitude of ways in which patient data can steer the risk assessment process.
For instance, employing an NL style, a total of $N$ few-shot examples like the following are generated (from training data), and integrated into the prompt template.
Few-shot Examples: Example 1.
Consider a patient with atypical angina, a resting blood pressure of 135 mmHg, and a serum cholesterol level of 250 mg/dl. Despite a fasting blood sugar below 120 mg/dl and normal resting electrocardiographic results, the patient exhibits a maximum heart rate of 160 bpm without exercise-induced angina. However, an ST depression of 1.5 is observed, with a slope of the peak exercise ST segment at 2, no major vessels colored by fluoroscopy, …
The risk of heart disease was assessed as high risk (’1’).
Example 2.
In contrast, within a numerical model, these examples would be presented as follows:
Few-shot Examples:
Example 1.
    age: 50.0
    sex: 1
    cp: 2.0
    trestbps: 170.0
    chol: 209.0
    fbs: 0
    restecg: 1.0
    thalach: 116.0
    exang: 0
    oldpeak: 0.0
    slope: 2.0
    0.0
    thal: 6.0
The risk of heart disease was assessed as high risk (’0’).
Example 2.
4.1.4. Infusion of data into a structured multi-layer prompt template
The core step in the prompt construction layer is the creation of a structured, multi-layer prompt template designed to facilitate the interpretation and analysis by the LLM, as illustaed in Figure 1. This involves the systematic arrangement of the user profile, domain knowledge, and few-shot examples into a format that is both logically coherent and tailored to the processing capabilities of LLMs.
  1. (1)
    Prompt Overview. At the beginning of the prompt, an overview section provides a concise explanation of the task context and objectives, ensuring that the LLM understands the broader goal of the clinical DM task it is about to perform.
  2. (2)
    Few-Shot Examples. The template incorporates few-shot examples to provide context and exemplify the application of domain knowledge to patient data, aiding the LLM in understanding the application of knowledge to new cases. The few-shot examples can be generated in both numerical and natural language modes, depending on the tested setting.
  3. (3)
    Target User Profile. The user profile generated in the previous steps is then integrated, transforming raw patient data into a textual narrative or maintaining a numerical representation, as required by the task design.
  4. (4)
    Task Description. A detailed task description is included, which clearly instructs the LLM on what it needs to do with the given information—such as predicting the risk of heart disease.
  5. (5)
    Integrating Domain Knowledge. Following the overview, the prompt incorporates domain knowledge using feature importance analysis. This section informs the LLM of the significant variables that should be considered when making the assessment, further grounding its predictions in the clinical knowledge base.
This multi-layer prompt template is an important component of the XAI4LLM framework, encapsulating the transformation of numerical and domain-specific data into a format optimized for the text processing capabilities of LLMs. It represents the final step in preparing the data for interaction with the LLM, ensuring that the model has all necessary context to make an informed and accurate prediction.
4.2. LLM Interaction and Prediction Layer
Having meticulously constructed a multi-layered prompt through the aforementioned Prompt Construction Layer, we transition into the LLM Interaction and Prediction Layer. This crucial phase embodies the application of the carefully tailored prompts to facilitate a meaningful interaction between the LLM and the clinical DM process. The primary aim is to harness the natural language processing abilities of LLMs to not only understand and analyze the patient data but also to generate informed and accurate predictions.
4.2.1. Communcation styles with LLMs
The Interaction and Prediction Layer navigates through two primary communication styles that an LLM might adopt in a clinical setting: Numerical Conversational (NC) and Natural-Language Single-Turn (NL-St). Both styles leverage the context and insights provided by the constructed prompts to enable the LLM to perform a virtual equivalent of a clinical evaluation. Whether through a step-by-step conversational approach or a comprehensive single narrative assessment, the LLM is tasked with distilling the complexities of the data into a definitive clinical prediction.
The success of this interaction layer depends upon the synergy between the informational structure of the prompts and the interpretative algorithms of LLMs. In the following sections, we will explore these communication styles in depth, elucidate the methods by which the LLM processes and responds to the prompts, and assess the efficacy of each approach in the context of heart disease risk prediction.
Example of Communication Styles in LLM Interaction - P1 Numerical Conversational (NC). Patient Interaction: The clinician conducts a step-by-step interview, gathering details such as age and sex, and continues to systematically collect all necessary patient information. Example.
 <User> Age: 54 years old
 <User> Sex: 1
 <User> Chest Pain Type: 3
 <User> Resting Blood Pressure: 150 mmHg
 <User> Serum Cholesterol: 223 mg/dl
LLM Response: Based on the sequentially provided data, the LLM evaluates the heart disease risk, emulating a methodical assessment of a clinician during an examination.
Example of Communication Styles in LLM Interaction - P2 Natural Language Single-Turn (NL-St). Complete Case Summary: A consolidated narrative provides a full overview of the health status of patients, incorporating data points like age, sex, chest pain type, blood pressure, cholesterol, and fasting blood sugar into one descriptive prompt. Example.
A 54-year-old man with a history of non-anginal chest pain, a resting
blood pressure of 150 mmHg, and a serum cholesterol level of 223 mg/dl
presents for evaluation ...
LLM Assessment: With the entire patient profile at hand, the LLM performs a risk assessment in one go, similar to a clinician evaluating a comprehensive medical report.
4.2.2. LLM Reaosoning
With the prompts ready, the LLM, enhanced with medical expertise, processes the information using one of two reasoning styles:
  1. (1)
    Direct Answer Extraction: The LLM provides a straightforward prediction based on the input data.
  2. (2)
    Chain-of-thought Reasoning (COT): Here, the LLM emulates a clinician’s thought process, narrating the reasoning behind its assessment.
The processing outcome is then refined through an NLP post-processing layer, culminating in a binary outcome that signifies the heart disease risk level, which is either ‘0’ for low risk or ‘1’ for high risk.
4.3. NLP Post-Processing
The responses of LLMs can sometimes be verbose, containing extensive text that encapsulates the required information within a comprehensive narrative. In particular, methods like Chain of Thought (COT) can generate even longer responses, embedding the conclusion in a detailed explanation. The primary challenge lies in distilling this abundant information into a binary outcome (0 or 1), which is crucial for applications that depend on straightforward, binary answers. To address this, our approach involves a two-tiered NLP post-processing strategy tailored to refine and extract the decisive answer from the LLM output.
Firstly, we guide the LMM to provide responses that conclude with a binary outcome (both in the prompt and system Role), aiming to encapsulate the summary information in the final sentence. However, LLMs may not always deliver a clear binary response. To tackle such instances, our NLP module employs a dual-layered approach: (i) An NLP filter that identifies and interprets specific keywords—out of a predefined set of 30 tailored to positive and negative outcomes—through different libraries to measure semantic proximity. (ii) A sentiment analysis tool activated if the keyword-based layer does not yield a conclusive outcome, further scrutinizing the text to derive the final binary decision.
This NLP Post-Processing mechanism plays a pivotal role in our, ensuring that the nuanced and sometimes verbose assessments provided by LMs are accurately transformed into binary classifications. This process respects the detailed nature of LM outputs and leverages linguistic cues and sentiment analysis to support precise and clinically relevant decision-making.
5. Experimental Configuration
Objective. The focus of this study is on binary classification tasks within the context of health risk evaluations using machine learning (ML) techniques. Specifically, it aims to learn a mapping function $f:\mathcal{X}\rightarrow\{0,1\}$ to predict a binary outcome $y\in\{0,1\}$ for each input feature instance $x\in\mathcal{X}$.
Data Collection. The Heart Disease dataset (Janosi et al., 1988), utilized in this study, was gathered from four medical institutions across the USA, Switzerland, and Hungary. The primary goal with this dataset is to perform binary classification to ascertain the presence or absence of heart disease. Initially, the dataset contained 621 out of 920 samples with missing entries, which were comprehensively imputed using the KNN technique, leading to a dataset devoid of missing values. The dataset exhibits a distribution of approximately 78% males to 21% females, and within this population, the prevalence of heart disease is higher among males (63%) compared to females (28%).
Benchmark Prediction Models. To evaluate our model performance against various degrees of data imbalance, we established three benchmark prediction models, each functioning independently of the input features. These models are: Maj Class-1 Prediction, which makes predictions with 90% of the outcomes being class 1, selected randomly; Maj Class-0 Prediction, that predicts with 90% of the outcomes as class 0, chosen at random; and Random Prediction, which makes its predictions entirely at random.
Model Hyperparameters and ML Algorithms. In our approach, we employed four distinct ML algorithms, each associated with its own set of hyperparameters. These parameters were optimized through a randomized search cross-validation (CV) approach, encompassing 29 unique hyperparameters across all algorithms. This optimization process involved a thorough tuning phase with numerous iterations of the models. We utilized a 5-fold CV scheme (0.9, 0.1) with RandomizedSearchCV across 20 iterations. The specific hyperparameters for each model are provided in Table 2.
Table 2. Summary of ML models and their Hyper-parameters
Algorithm Parameters Total
RF n_estimators, max_depth, min_samples_split, min_samples_leaf, bootstrap 5
MLP hidden_layer_sizes, activation, solver, alpha, learning_rate, learning_rate_init, tol, max_iter 8
XGB use_label_encoder, eval_metric, n_estimators, learning_rate, max_depth, colsample_bytree 6
LightGBM n_estimators, learning_rate, max_depth, num_leaves, min_data_in_leaf, feature_fraction, bagging_fraction, bagging_freq 24
API Model of OpenAI. For assessing the risk of heart disease from patient health records, we integrated OpenAI GPT-3.5-turbo model. This model facilitates interaction through a conversational interface, processing a sequence of messages and user inputs as prompts to produce relevant responses. The model temperature was set to 0, ensuring deterministic output behavior. Utilization of the OpenAI model included batching requests containing patient data features, with each batch processed to yield a prediction. The implementation leveraged the OpenAI API, the tenacity library for retrying upon failure, along with standard data processing libraries. For replication purposes, the source code has been made accessible. [^5] Source code available at https://l1nk.dev/EbrXw
Categorization Process. Our domain knowledge enrichment module has been designed to meticulously categorize features into levels of significance by analyzing their importance scores. This process begins by separating features into positive and negative impacts, with further division into high, medium, and low importance based on quantile thresholds. Specifically, features are evaluated by calculting the 33rd and 67th quantiles for both positive and negative importance scores, establishing thresholds for categorization.
Features with importance scores above the 67th quantile threshold for positive impacts are deemed of high importance, while those between the 33rd and 67th quantiles are considered of medium importance. Similar logic applies to features with negative impacts, categorizing them based on their deviation from neutral. Readers are referred to Figure 3 for an more in-depeth understanding.
Visualization and Analysis To facilitate a deeper understanding and enable intuitive analysis, we visualize the categorized features using a scatter plot that reflects their importance scores and categorizations. The plot incorporates color coding to differentiate between the categories and includes lines marking the quantile thresholds. This visualization provides a clear representation of the impact of each feature, offering insights into how different indicators influence the decision-making process of the model. By employing this structured approach to feature categorization, we enhance the ability of LLM to prioritize information based on its relevance to clinical outcomes, ensuring a more informed and accurate assessment process.
In our experimental framework, we operationalize the evaluation of model performance through a series of metrics that emphasize not only prediction accuracy but also the cost-sensitivity aspect of the predictions. These metrics are designed to reflect the effectiveness of the model in a clinical decision-making context.
5.1. Key Evaluation Metrics
Two categories of evaluation metrics are utilized: accuracy and risk, each addressing the positive and negative aspects of LLMs. For accuracy, we primarily employ the F1 score to conduct our analysis. In assessing risk, we focus on false negatives—due to their significant impact—and gender bias, identifying these as two distinct categories of risks.
Accuracy and Risk Sensitivity
We report on all metrics, including accuracy, precision, recall, and the $F_{\beta}$ score. Our analysis primarily utilizes the $F_{\beta}$ score, which generalizes the F1 score by allowing us to adjust the importance of precision versus recall through the $\beta$ parameter. The formula for calculating the $F_{\beta}$ score is:
(6) $F_{\beta}=(1+\beta^{2})\times\frac{\text{Precision}\times\text{Recall}}{\beta^{2}\times\text{Precision}+\text{Recall}}$
where Precision is $\frac{TP}{TP+FP}$, Recall is $\frac{TP}{TP+FN}$, and $TP$, $FP$, and $FN$ represent the numbers of true positives, false positives, and false negatives, respectively.
In this study, we utilize two specific instances of the $F_{\beta}$ score to cater to different aspects of clinical decision-making:
  1. (1)
    For General Performance Evaluation: We use $\beta=1$ (yielding the F1 score) to balance precision and recall. This choice reflects our interest in evaluating the performance of the model where both false positives and false negatives carry similar weight.
    (7) $F_{1}=2\times\frac{\text{Precision}\times\text{Recall}}{\text{Precision}+\text{Recall}}$
  2. (2)
    For Risk Sensitivity: Considering the critical nature of clinical outcomes, we also use $\beta=3$, yielding the F3 score. This score gives more weight to recall, reflecting the higher cost of false negatives in clinical settings.
    (8) $F_{3}=\frac{(1+3^{2})\times\text{Precision}\times\text{Recall}}{3^{2}\times\text{Precision}+\text{Recall}}$
By adopting the F3 score, we emphasize the importance of minimizing false negatives due to the grave consequences of overlooking at-risk individuals. This approach ensures that our model evaluation aligns with the prioritization of patient safety and effective risk identification.
Fairness Metrics
In assessing fairness, we focus on the following metrics which are crucial for ensuring equitable treatment across different demographic groups.
Equality of Opportunity ($\Delta_{EO}$) Equality of opportunity is assessed by the difference in true positive rates between groups $\Delta_{EO}=|\text{TPR}_{M}-\text{TPR}_{F}|$ with TPR representing the true positive rate. Predictive Parity ($\Delta_{PP}$) Predictive parity requires that different groups have equal positive predictive values: $\Delta_{PP}=|\text{PPV}_{M}-\text{PPV}_{F}|$, where PPV is the positive predictive value. Predictive Equality Accuracy ($\Delta_{PEAc}$). Finally, the difference in accuracy across genders is defined by $\Delta_{PEAc}=|\text{Acc}_{M}-\text{Acc}_{F}|$.
6. Results
6.1. Research Questions.
To gain a better understanding of the merits of the proposed system, we have conducted a series of experiments to address the research evaluation questions (RQs) listed below:
  • The accuracy and risks of false predictions in LLM-based Models in zero-shot (RQ1) and few-shot learning settings (RQ2).
  • The bias and risk of unfairness from a gender perspective (RQ3).
To achieve our objectives, we have formulated the following detailed research questions (RQs):
  • RQ1.
    In the zero-shot learning setting, how do LLM-based models perform compared to classical ML models?
  • RQ2.
    How do the results of zero-shot learning change when transitioning to a few-shot learning setting?
  • RQ3.
    What is the extent of bias and the risk of unfairness in LLM-based, when focusing on gender bias?
Key Catalysts of the Research Our investigation specifically aims to explore the empirical Research Questions (RQs) through the lens of key catalysts of this research. These catalysts serve as pivotal variables that influence the outcomes and insights derived from our study, delineated as follows: (1) Prompt Design Strategies Zero-shot learning – Evaluating the adaptability of the model to tasks without prior examples. Few-shot learning – Assessing the performance of the model with a minimal number of examples. (2) Integration of Domain Knowledge. Obtained by means of feature-level explanation driven from a suite of ML models, (a) RF (b) MLP (c) XGB (d) LightGBM (3) Narrative Style Direct narrative – Analyzing the direct question approach on model comprehension. Chain of Thought (COT) – Exploring the reasoning process and its effect on model outcome. (4) Communication Style Numerical Conversational (NC) style – Evaluating the impact of dialogue in model reasoning and prediction accuracy. Natural Language Single-Turn Profile (NL-St) – Gauging the efficacy of a cohesive narrative in eliciting accurate model responses.
In the following sections, we provide a detailed analysis and response to the aforementioned research questions.
Answers to RQ1. Given the detailed research questions (RQs) and the summary of results from the table provided, an analysis of the performance of LLM-based models in a zero-shot learning setting (RQ1) is structured around two key metrics: accuracy (assessed through F1 score) and risks (assessed through F3 score, which gives more attention to recall, thus cost-sensitive). The analysis focuses on the performance of models employing XAI as domain knowledge and compares them across different numerical communication styles (NC) and Natural-language Single Turn (NL-St) presented on the left and right side of Table 3, respectively. Note that, rows for LLM (above dashed line) refer to direct or COT reasoning style.
Table 3. Model Performance Metrics for NC Style and NL-Sp Style in Zero-shot ICL setting. Notable performances discussed in the body of the text are highlighted in bold.
Domain Knowledge NC Style NL-St Style
Model XAI-model Acc. Pre. Rec. $F_{\beta}$ Acc. Pre. Rec. $F_{\beta}$
$\beta=1$ $\beta=3$ $\beta=1$ $\beta=3$
Direct 0.576 0.583 0.839 0.687 0.802 0.660 0.665 0.784 0.719 0.770
COT 0.598 0.607 0.779 0.682 0.758 0.622 0.610 0.897 0.726 0.856
Direct RF 0.620 0.595 0.980 0.741 0.921 0.630 0.639 0.765 0.696 0.750
COT RF 0.576 0.597 0.725 0.655 0.710 0.620 0.614 0.843 0.711 0.813
Direct MLP 0.609 0.597 0.902 0.719 0.858 0.543 0.588 0.588 0.588 0.588
COT MLP 0.554 0.581 0.706 0.637 0.691 0.674 0.672 0.804 0.732 0.788
Direct XGB 0.576 0.583 0.824 0.683 0.791 0.728 0.760 0.745 0.752 0.747
COT XGB 0.576 0.591 0.765 0.667 0.743 0.707 0.700 0.824 0.757 0.809
Direct LightGBM 0.609 0.609 0.824 0.700 0.795 0.500 0.541 0.647 0.589 0.635
COT LightGBM 0.609 0.647 0.647 0.647 0.647 0.609 0.647 0.647 0.647 0.647
[…]
MLP 0.815 0.827 0.843 0.835 0.841 0.815 0.827 0.843 0.835 0.841
XGB 0.902 0.889 0.941 0.914 0.936 0.902 0.889 0.941 0.914 0.936
LightGBM 0.891 0.887 0.922 0.904 0.918 0.891 0.887 0.922 0.904 0.918
Random 0.500 0.547 0.569 0.558 0.566 0.500 0.547 0.569 0.558 0.566
Majority Zeros 0.402 0.300 0.059 0.098 0.064 0.402 0.300 0.059 0.098 0.064
Majority Ones 0.489 0.524 0.843 0.647 0.795 0.489 0.524 0.843 0.647 0.795
  • In the zero-shot ICL setting, the integration of XAI through domain knowledge such as RF, MLP, XGB, and LightGBM has shown varied effects on the performance of LLM-based models. Specifically, models integrating RF domain knowledge under NC style exhibited a notable increase in accuracy (F1: 0.741) and cost-sensitive accuracy (F3: 0.921), indicating a superior capability to minimize false predictions while maintaining high accuracy.
  • This contrasts with models in the NL-St style, which showed slightly lower performance metrics. Moreover, comparing narrative styles, the Chain of Thought (COT) approach, particularly with XGB domain knowledge under the NL-sT style, demonstrated substantial resilience in managing risks (F3: 0.809) while ensuring satisfactory accuracy (F1: 0.757). This suggests that the COT narrative, combined with XAI techniques, particularly benefits the NL-St communication style by effectively reducing the risk of false predictions, a critical factor in zero-shot learning scenarios where models operate without prior examples.
  • The analysis reveals that the integration of domain knowledge in LLM-based models significantly impacts their performance in zero-shot learning settings. Models utilizing RF and XGB domain knowledge, especially when employing a COT narrative in an NL-sp style, show promising results in balancing accuracy and risk management. These findings underscore the importance of selecting appropriate domain knowledge and narrative styles to optimize LLM-based model performance, addressing key challenges in zero-shot learning and paving the way for advancements in AI systems’ development and application.
  • While LLM-based models integrating XAI domain knowledge, such as the model employing RF with NC style (F1: 0.741), do not match the superior F1 scores of traditional ML models (e.g., XGB’s F1 of 0.914), their performance when evaluated with F3—a metric that significantly values minimizing false predictions—shows noteworthy strength (e.g., RF with NC style reaching an F3 score of 0.921). This illustrates that, despite the inherent limitation of operating without specific task-related information, LLM-based models equipped with XAI can achieve a reliable balance between accuracy and risk management. For instance, the COT narrative combined with XGB domain knowledge under the NL-St style yields an F3 of 0.809, underscoring these models’ potential in scenarios where reducing false predictions is paramount. This competency highlights the models’ viability as dependable resources alongside general GPT models in diverse settings.
Overall, the results of the first analysis provides insight into optimizing LLM-based models for accuracy and reducing false predictions, contributing to the broader research on ethical AI development by highlighting the effective integration of XAI and the strategic use of narrative vs. conversational styles.
Answers to RQ2. We can analyze the shift from zero-shot to few-shot learning in terms of F1 and F3 scores in Figure 4, reflecting standard accuracy and risk-aware accuracy respectively.
(a) NC style
(b) NL-St style
Figure 4. Comparison of performance in the few-shot setting and across two communication styles with LLMs.
  • In the NC style (top graphs), there is a discernible improvement in F1 scores for both Direct and COT narratives as the number of shots increases, with the models as the number of shots increases, with the models stabilizing after a certain point. This trend suggests that few-shot learning enhances the accuracy of the models as they benefit from the additional context provided by the examples. The F3 scores in the NC style also generally improve with the number of shots, although the pattern is less consistent, reflecting the models’ varying sensitivity to the cost of false predictions. Notably, the models incorporating COT narratives tend to maintain or achieve higher F3 scores, especially with RF and XGB, indicating a more robust risk management strategy in the few-shot learning setting.
  • In contrast, the NL-St style (bottom graphs) reveals a more volatile trajectory for F1 scores, particularly for the COT narratives, which initially see a sharp increase in accuracy but then experience fluctuations. This could suggest that while the COT narrative helps the model understand the context with a few examples, too many examples might introduce complexity or noise. The F3 scores for the NL-St style, especially for the COT narratives, show a peak at a moderate number of shots before declining, underscoring the importance of an optimal number of examples for managing risks effectively.
  • Across both communication styles, the F1 and F3 scores show that while models generally benefit from few-shot learning, there is a nuanced interplay between the number of shots and the narrative style that affects their performance. The models perform better in terms of risk-aware accuracy with a balanced number of shots and appear to leverage the structured reasoning provided by the COT narrative to manage false predictions better.
Overall, the transition to few-shot learning settings illustrates that LLM-based models not only improve in conventional accuracy (F1) with the right amount of training data but also enhance their ability to manage risks (F3) effectively. The data implies an optimal range of examples that maximize model performance without introducing additional risks, which is particularly crucial for applications where the cost of errors is high. The findings also highlight the significance of choosing the appropriate narrative and communication style to tailor the models’ learning process to specific tasks and contexts.
Table 4. Fairness Evaluation for Gender Groups (Top ZeroShot, Bottom Few Shot with 16 examples). Results only focus on the NC style for simplification.
Model DW RecM RecF $\Delta_{EO}$ a PrecM PrecF $\Delta_{PP}$ AccM AccF $\Delta_{PEAc}$
Direct RF 0.980 1.000 0.020 0.676 0.154 0.522 0.684 0.312 0.372
COT RF 0.714 1.000 0.286 0.686 0.182 0.504 0.605 0.438 0.167
Direct MLP 0.918 0.500 0.418 0.682 0.091 0.591 0.671 0.312 0.359
COT MLP 0.714 0.500 0.214 0.700 0.083 0.617 0.618 0.250 0.368
Direct XGB 0.816 1.000 0.184 0.656 0.182 0.474 0.605 0.438 0.167
COT XGB 0.755 1.000 0.245 0.673 0.182 0.491 0.605 0.438 0.167
Direct LightGBM 0.816 1.000 0.184 0.702 0.167 0.535 0.658 0.375 0.283
COT LightGBM 0.633 1.000 0.367 0.721 0.250 0.471 0.605 0.625 0.020
[…]
MLP - 0.837 1.000 0.163 0.854 0.500 0.354 0.803 0.875 0.072
XGB - 0.939 1.000 0.061 0.920 0.500 0.420 0.908 0.875 0.033
LightGBM - 0.918 1.000 0.082 0.918 0.500 0.418 0.895 0.875 0.020
Direct RF 0.939 1.000 0.061 0.730 0.222 0.508 0.737 0.562 0.175
COT RF 0.857 1.000 0.143 0.667 0.286 0.381 0.632 0.688 0.056
Direct MLP 1.000 1.000 0.000 0.690 0.200 0.490 0.711 0.500 0.211
COT MLP 0.755 0.500 0.255 0.685 0.167 0.518 0.618 0.625 0.007
Direct XGB 1.000 1.000 0.000 0.700 0.182 0.518 0.724 0.438 0.286
COT XGB 0.980 1.000 0.020 0.676 0.182 0.494 0.684 0.438 0.246
Direct LightGBM 0.918 1.000 0.082 0.750 0.200 0.550 0.750 0.500 0.250
COT LightGBM 0.878 1.000 0.122 0.729 0.222 0.507 0.711 0.562 0.149
Answer to RQ3. Table 4 evaluates the fairness of various ML models from a gender perspective, focusing on RF, MLP, XGB, and LightGBM. Fairness metrics such as recall, precision, and accuracy differences between male (M) and female (F) gender groups ($\text{Rec}_{M},\text{Rec}_{F},\text{Prec}_{M},\text{Prec}_{F},\text{Acc}_{M},\text{Acc}_{F}$) and biases in equality of opportunity ($\Delta_{EO}$), predictive parity ($\Delta_{PP}$), and predictive equality accuracy ($\Delta_{PEAc}$) are assessed (Verma and Rubin, 2018). Note that here, rows for LLM refer to direct or COT reasoning style, and we only present results of gender fairness for NC communication style due to its superior performance in the previous sections.
Traditional ML models without LLM-based strategies show relatively small disparities in recall, with $\Delta_{EO}$ suggesting lower bias in identifying true cases across genders. However, a notable difference in precision ($\Delta_{PP}$) is observed, particularly with RF and XGB models, where males are predicted more precisely than females. This discrepancy is somewhat mitigated when certain LLM strategies are integrated (e.g., RF using COT with n=16). However, the overall results regarding precision and fairness are comparable, or at least not significantly different, between classical and LLM-based models.
The analysis of accuracy differences ($\Delta_{PEAc}$) reveals that traditional ML models maintain minimal gender bias, with disparities ranging from 0.007 to 0.072. In contrast, when applying Direct in conjunction with domain knowledge, the difference in accuracy escalates, as seen with the Direct RF model ($\Delta_{PEAc}=0.372$). Conversely, employing COT narratives appears to significantly reduce this bias (e.g., COT with LightGBM yields $\Delta_{PEAc}=0.020$), suggesting the effectiveness of COT in fostering fairness.
In conclusion, the overall trend is that traditional ML models tend to exhibit lower bias in recall and accuracy. However, there are notable cases where the integration of certain reasoning strategies, few-shot learning, and XAI models reduces the bias even further than ML cases (e.g., with n=16, both direct and COT produce with XGB the lowest recall equal to 0.000 and 0.0020 respectively). In other words, the application of LLM strategies, especially COT narratives, seems to enhance predictive parity and equality in accuracy between genders. The result of gender bias appears to be less evident in precision bias between ML and LLM-based techniques. This underscores the ongoing necessity to develop algorithms that achieve both high accuracy and equitable treatment across diverse demographic groups. Surprisingly, despite the general trend of higher bias observed in LLM applications compared to traditional ML models, LLMs demonstrate an unexpectedly higher recall for the underrepresented gender group (females). This trend suggests that while LLMs might introduce certain biases, they might have the advantage in recognizing cases among traditionally underrepresented groups, potentially mitigating historical disparities in healthcare outcomes.
7. Conclusion
This research focuses on integrating large language models (LLMs) into healthcare decision-making, especially for heart risk prediction through zero/few-shot in-context learning (ICL). The primary contribution of this work is the development of a systematic structured prompt construction that has been shown enhancing diagnostic accuracy of LLMs by incorporating domain-specific medical knowledge into their decision-making process. This includes creating multi-layered structured templates that efficiently integrate numerical and natural language user profiles along with critical medical domain knowledge. Furthermore, this study enhances our understanding of the optimal communication styles with LLMs—whether conversational or single-long prompt—that can best utilize LLMs in clinical diagnostics.
To provide a clearer summary of the results and key takeaways, we present the main findings below and in Figure 5, which includes a checklist of the primary results obtained.
The results of the study are promising, demonstrating that the structured approach to prompt construction and the nuanced integration of communication styles can significantly impact the diagnostic precision of LLMs in healthcare. By comparing these new ways of prediction compared to traditional ML models, the research highlights the potential of LLMs to revolutionize medical diagnostics by offering more adaptive, nuanced, and potentially resource-conserving methods. Additionally, it emphasizes the importance of addressing inherent risks such as false negatives and biases, which are critical for ensuring the safe and equitable application of LLMs in clinical settings. Through these insights and methodologies, this research paves the way for more effective and nuanced applications of AI in healthcare.
Figure 5. Illustration of Key Findings. The x-axis represents the number of few-shot examples, directly correlating to the number of communication turns in the Numerical Conversational (NC) style and influencing the content of the Natural Language (NL) profile. The y-axis denotes both precision and risk-sensitive precision, measured by F1 and F3 scores, showcasing the performance of our LLMs in medical diagnostic tasks.
Summary of LLM Diagnostic Performance Analysis Main Findings: The graph in Figure 5 illustrates the relationship between the length of communication or prompt (synonymous with the number of few-shot examples, ranging from 0 to 32) and the resulting diagnostic performance measured by an accuracy/cost-sensitive accuracy metric. (1) Conversational Style Achieves Higher Performance. The “Numerical Conversational (NC)” style shows improvement with an increasing number of few-shot examples and surpassed the “Natural language Single Turn (NL-St)” prompt style in achieving peak accuracy, indicating that conversational interaction is potentially more effective for long LLM diagnostic description. (2) CoT Reasoning Aids Textual Presentation. Chain-of-thought reasoning maintains quality performance with growing example counts, within the narrative single-prompt context. This advantage points to effectiveness of COT in helping LLMs manage and utilize larger amounts of textual data without a decline in diagnostic accuracy. Surprisingly, COT has an adverse impact in the conversational mode. (3) The Adverse Impact of Long Prompts on Quality of Predictions. In contrast to the conversational model, which appears less sensitive to the number of examples, the quality of the NL-St profile declines with increasing prompt length. This suggests that the conversational mode is better suited to digest a larger volume of examples, processing them sequentially, one by one, and may imply that it is better equipped to manage and interpret extensive data sets.